a wire 80 cm long has a mass of 10 g. find the velocity of a transverse wave in the wire if it is under a tension of 9.0x10 to 11th dynes.


1
Expert's answer
2020-10-15T10:45:59-0400
"v=\sqrt{\frac{Tl}{m}}\\v=\sqrt{\frac{9\cdot10^{11}(10^{-5})\cdot 0.8}{0.01}}\\v=27000\frac{m}{s}"